# IEEEexample.bib

### v2

@@ Abstract
Methane ($CH_{4}$) is the second most prevalent greenhouse gas after carbon dioxide ($CO_{2}$), with a short atmospheric lifetime of about 9.14 years. Reducing methane emissions could quickly mitigate global warming. A methane plume is a localized methane concentration. Methane emission monitoring involves three tasks: concentration inversion, plume segmentation, and emission rate estimation. The matched filter algorithm (e.g., Mag1c) for methane concentration inversion is sensitive to noise and reference signals. Research on methane plume segmentation is limited, often relying on subjective manual segmentation. Emission rate estimation typically uses the integrated mass enhancement (IME) algorithm, which depends on accurate meteorological measurements, limiting its applicability. Using PRISMA hyperspectral images from the WENT landfill in Hong Kong and the EnMAP hyperspectral dataset as a base map, we propose a deep learning framework for quantitative methane emission monitoring from hyperspectral images based on physical simulation. We simulate methane plumes using large eddy simulation (LES) and create various fugitive methane concentration maps using the radiative transfer equation (RTE). We further augment data to create a simulated EnMAP dataset $(https://github.com/Sigx093105/CH4_{p}lume_{d}ataset.git)$. We train a U-Net for concentration inversion, a Mask R-CNN for plume segmentation, and a ResNet-50 for emission rate estimation. These deep networks outperform conventional algorithms in validation accuracy and generalize well on RPISAM data. We combine tasks into multi-task learning models: MTL-01 and MTL-02, which outperform single-task models. This research demonstrates multi-task deep learning for quantitative methane monitoring and can be extended to various methane monitoring tasks.
@@ I Introduction
To detect emitted methane plumes, airborne data (AVIRIS-NG) is commonly utilized due to small-scale methane emission sources [foote2020fast][foote2021impact]. However, the advent of PRISMA, a new-generation hyperspectral satellite with a spatial resolution of 30m, has now made it feasible to retrieve spaceborne methane emissions [cogliati2021prisma]. The monitoring process for methane plumes can be divided into three tasks: (1) methane concentration inversion, (2) methane plume segmentation, and (3) estimation of single plume flux rates. For methane concentration inversion, current methods predominantly rely on differential optical absorption spectroscopy (DOAS) and matched filter techniques. DOAS, although computationally intensive, is typically used for small-scale inversion in scenarios where emission point sources are known [thorpe2017airborne][thorpe2013high]. On the other hand, matched filter algorithms are commonly applied for large-scale methane plume screening [foote2021impact][thompson2015atmospheric]. However, it is important to note that the accuracy of the matched filter algorithm is heavily based on the precise modeling of the target spectrum and background estimation [foote2021impact][theiler2006effect01].
@@
Few studies have been conducted on methane plume segmentation, but it is crucial to accurately delineate single plumes, as it significantly determines the estimation accuracy of methane flux rate. Over the past few decades, a plethora of algorithms have been proposed for estimating point source emission rates. However, most of these algorithms require local wind speed data as auxiliary information, which is difficult to obtain due to the limited availability of meteorological observation stations[varon2018quantifying] [jongaramrungruang2022methanet]. addressed this limitation by employing simulated plumes and deep learning techniques to estimate emission rates without the need for wind speed information, although the presented network could only estimate the emission rate of single-source plumes. In practical scenarios, it is often necessary to segment individual plumes from an image and estimate the emission rates separately.
@@
Since the introduction of AlexNet in 2012, deep learning has witnessed rapid development and yielded remarkable results in various fields of remote sensing application[JIA202314][JIANG202330][he2024ubfineneturbanbuildingfinegrained]. The U-net deep learning architecture was initially designed for semantic segmentation tasks[ronneberger2015u][POLEWSKI2021297], but also exhibits enormous potential for regression[10.1145/3429309.3429328]. We regard methane concentration inversion as pixel-by-pixel regression, so U-net can be used to complete the task. Additionally, Mask R-CNN has proven to be highly effective in a wide range of instance segmentation tasks [SANIMOHAMMED2022100024] and should also be applicable to Methane plume segmentation. Moreover, ResNet addresses the issue of gradient vanishing, enabling deeper networks to improve feature extraction capabilities[he2016deep][shafiq2022deep]. Many of the former works for emission rate estimation using shallow convolutional neural networks(CNN) have proven successfully to be realizable, but with moderate precision[9027926][guanter2021mapping]. It is reasonable to apply ResNet to improve the emission rate estimation. And our goal is to leverage physically simulated plumes as training data for multi-task deep learning techniques to simultaneously perform methane concentration inversion, methane plume segmentation, and flux rate estimation in a unified approach.
@@
In this article, we start by describing the foundational principles of the algorithms utilized in our study, including large-eddy simulation, radiation simulation, the methane plume concentration inversion algorithm, the methane plume segmentation algorithm (instance segmentation), and the plume emission rate estimation algorithm. Following this, we provide a detailed account of our experimental setup, which encompasses the research area, the collected data, and the procedures used to build our methane monitoring dataset. Lastly, we evaluate the performance of deep learning algorithms against traditional ones for the tasks of methane plume concentration inversion, methane plume segmentation (instance segmentation), and plume emission rate estimation. Furthermore, we present the two multi-task learning networks, MTL-01 and MTL-02, that we developed for these tasks. The major contributions of this paper are as follows:
@@
1. We propose a physics-informed deep learning scheme for spaceborne retrieval of methane emissions based on data simulation using Large Eddy Simulation (LES), Radiative Transfer Equation (RTE), and data augmentation techniques.
@@
2. We apply an instance segmentation algorithm to solve the problem of isolating and identifying methane emission sources by plume segmentation.
@@
3. We show that the serialization of multiple sub-tasks for methane emission monitoring will lead to additional error accumulation, and designed a multi-task learning model to suppress the errors.
@@ II-A Simulation
The simulation operation consists of two parts: simulation of methane plume emission and simulation of EnMAP-like satellite images with methane plume signal. For methane plume simulation, we utilized Large Eddy Simulation (LES), which is a computational fluid dynamics (CFD) method used to simulate turbulent phenomena in fluid flow[sagaut2002large]. It combines the advantages of Direct Numerical Simulation (DNS) and Reynolds-averaged Navier-Stokes equations (RANS) methods, aiming to accurately predict the statistical characteristics of turbulent flow[pope2001turbulent][domaradzki1997subgrid]. The core of LES lies in filtering the flow field into large- and small-scale vortices through a filtering operation. This can be achieved by applying a low-pass filter to remove high-frequency small-scale vortices, while retaining the low-frequency large-scale vortices[domaradzki1997subgrid]. The filtered flow field equations form the basis of LES. The governing equations for LES are based on Reynolds-averaged Navier-Stokes equations (RANS)[germano1991dynamic], which are solved to simulate motion and turbulent effects in the fluid [10.1115/1.3124648]:
@@
By combining these two steps, we can simulate methane plumes of various shapes, sizes, and concentrations and inject them into EnMAP data to generate simulated hyperspectral satellite images, which contain methane plume signal overlaid with a realistic scene background. This provides a solid foundation for our research and allows us to validate the performance and reliability of our methodology in real-world scenarios.
@@ II-B Methane Concentration Inversion
Methane concentration inversion in the Western New Territories (WENT) region was performed using PRISMA data and the mag1c method. The results of this inversion were evaluated against a deep learning model to validate its generalization capabilities. Considering the intricate distribution of land and sea areas in Hong Kong, as well as the challenging rugged terrain and dense urban construction, it can be expected that the surface conditions in the region significantly influence the results of methane concentration inversion. To address this, we utilize a K-means-based masking method to run Mag1c[funk2001clustering].
The U-net is a deep learning network characterized by an encoding-decoding structure and is primarily used for semantic segmentation tasks[ronneberger2015u][POLEWSKI2021297][zang2021land]. Typically, the input and output channels of the U-net are of equal size. However, to accommodate the requirements for the inversion of methane concentration, an additional layer of $1*1$ convolution is added after the U-net encoding-decoding structure[lin2013network], and the loss function is replaced for the regression task. The loss functions used include the mean squared error (MSE) and $SmoothL_{1}$:
@@
Prior to starting the training phase, various preprocessing techniques were applied to both the labels and images. In particular, we computed the mean and standard deviation of all non-zero pixel values in the labels to achieve normalization. Referring to the preprocessing methods of the DOAS algorithm, the subsequent formula was utilized to preprocess the images[wagner2004max][platt2008differential]:
@@
where $I_{before}$ refers to the preprocessed image, and $I_{after}$ represents the same image after preprocessing. Both EnMAP and PRISMA exhibit pixel values within the range of 0-10 after applying the logarithm, and dividing these values by 10 ensures that all preprocessed pixel values fall within the range of 0-1. This approach facilitates the generalization of the model across different sensors. The network is trained using two loss functions, MSE and SmoothL1, with an initial learning rate of 0.001. The learning rate is reduced by 75% every 10 epochs, totaling 200 epochs.
@@ II-C Methane Plume Segmentation
Instance segmentation is a classical task in computer vision, and methane plume segmentation is essentially an instance segmentation task. Previous research work on methane plume segmentation is limited, and many studies are based on subjective segmentation criteria[gorrono2023understanding]. Some studies suggested using traditional algorithms for the segmentation of methane plumes, with contour tracking algorithms considered a good choice due to their efficiency and stability. On the other hand, there are many deep learning-based instance segmentation algorithms, among which Mask R-CNN is the most classical and widespread one[he2017mask]. It is a two-step instance segmentation algorithm that has been proven to achieve high performance in a large number of datasets for a wide range of applications. Although new instance segmentation algorithms are continuing to emerge, some of the algorithms sacrifice segmentation accuracy for real-time segmentation[terven2023comprehensive], which is not necessarily required for our case. Other algorithms may achieve higher instance segmentation accuracy[wang2020solov2], but require increasingly complex optimization processes during training, which is not cost-effective for our case. Therefore, we chose Mask R-CNN as the deep learning algorithm for methane plume segmentation and active contour algorithms as the traditional counterpart algorithm. Typically, Mask R-CNN takes three image channels as inputs, but we modified the input layer to allow flexible alteration between 1 or 41 input channels ( the methane distribution map has only one band, and EnMAP remains 41 bands after excluding water vapor and carbon dioxide sensitive bands[foote2020fast]) to accommodate different segmentation requirements. We used the original loss function of Mask R-CNN:
@@
where $L$ is the total loss function, $L_{cls}$ is the category loss function, which is based on the cross-entropy loss, $L_{box}$ is the bounding box loss function, which is based on the $SmoothL_{1}$ loss, and $L_{mask}$ is the mask loss function, which typically uses the binary cross entropy:
@@
$L_{box}=-\text{IoU}(b,\hat{b})+\frac{\rho^{2}(b,\hat{b})}{c^{2}}$ (10)
where $b$ and $\hat{b}$ respectively denote the ground truth bounding box and the predicted bounding box, $\text{IoU}(b,\hat{b})$ indicates the intersection over the union between these two bounding boxes, $\rho^{2}(b,\hat{b})$ is the Euclidean distance between the center points of two bounding boxes, and $c$ is a hyperparameter typically set to 10.
II-D Methane Emission Rate Estimation
There are two scenarios for estimating emission rates. The first scenario involves estimating the emission rates using training samples specifically collected for this purpose. The goal here is to evaluate the performance of the estimation algorithm under ideal conditions. The second scenario refers to the estimation of emission rates based on plume segmentation results (MTL-01), where the objective is to evaluate and improve the accuracy of the plume instance segmentation algorithms.
@@ II-D Methane Emission Rate Estimation
ResNet is a deep learning method that uses residual blocks, which enables the neural network to learn the residual map by effectively avoiding issues such as gradient vanishing and explosion[he2016deep]. ResNet is widely applied in various application fields, including image recognition, object detection, and speech processing, yielding impressive results[veit2016residual][shafiq2022deep]. We trained the ResNet-50 network on samples for emission rate to perform the flux rate estimation task. Subsequently, ResNet-50 was used as a pretrained network to validate the rationale of plume segmentation results and assist in optimizing the instance segmentation network.

### v3

@@ Abstract
Methane ($CH_{4}$) is the second most prevalent greenhouse gas after carbon dioxide ($CO_{2}$), with a short atmospheric lifetime of approximately 9.14 years. Reducing methane emissions could quickly mitigate global warming. A methane plume is a localized concentration of methane. Methane emission monitoring involves three tasks: concentration inversion, plume segmentation, and emission rate estimation. The matched filter algorithm (e.g., Mag1c) for methane concentration inversion is sensitive to noise and reference signals. Research on methane plume segmentation is limited, often relying on subjective manual segmentation. Emission rate estimation typically uses the integrated mass enhancement (IME) algorithm, which is dependent on accurate meteorological measurements, limiting its applicability. Using PRISMA hyperspectral images from the WENT landfill in Hong Kong and the EnMAP hyperspectral dataset as a base map, a deep learning framework is proposed for quantitative methane emission monitoring from hyperspectral images based on physical simulation. We simulate methane plumes using large eddy simulation (LES) and create various fugitive methane concentration maps using the radiative transfer equation (RTE). We further augment the data to create a simulated EnMAP dataset $(https://github.com/Sigx093105/CH4_{p}lume_{d}ataset.git)$. We train a U-Net for concentration inversion, a Mask R-CNN for plume segmentation, and a ResNet-50 for emission rate estimation. These deep networks outperform conventional algorithms in validation accuracy and generalize well to RPISAM data. We combine tasks into multi-task learning models: MTL-01 and MTL-02, which outperform single-task models. This research demonstrates multi-task deep learning for quantitative methane monitoring and can be extended to various methane monitoring tasks.
@@ I Introduction
To detect emitted methane plumes, airborne data (AVIRIS-NG) is commonly utilized due to small-scale methane emission sources [foote2020fast][foote2021impact]. However, the advent of PRISMA, a new-generation hyperspectral satellite with a spatial resolution of 30 m, has made it feasible to retrieve spaceborne methane emissions [cogliati2021prisma]. The monitoring process for methane plumes can be divided into three tasks: (1) methane concentration inversion, (2) methane plume segmentation, and (3) estimation of single plume flux rates. For methane concentration inversion, current methods predominantly rely on differential optical absorption spectroscopy (DOAS) and matched filter techniques. DOAS, although computationally intensive, is typically used for small-scale inversion in scenarios where emission point sources are known [thorpe2017airborne][thorpe2013high]. On the other hand, matched filter algorithms are commonly applied for large-scale methane plume screening [foote2021impact][thompson2015atmospheric]. However, it is important to note that the accuracy of the matched filter algorithm is heavily based on the precise modeling of the target spectrum and background estimation [foote2021impact][theiler2006effect01].
@@
Few studies have been conducted on methane plume segmentation, but it is crucial to accurately delineate single plumes, as it significantly determines the accuracy of the estimation of the methane flux rate. Over the past few decades, a plethora of algorithms have been proposed for estimating point-source emission rates. However, most of these algorithms require local wind speed data as auxiliary information, which is difficult to obtain due to the limited availability of meteorological observation stations[varon2018quantifying] [jongaramrungruang2022methanet]. addressed this limitation by employing simulated plumes and deep learning techniques to estimate emission rates without the need for wind speed information, although the presented network could only estimate the emission rate of single-source plumes. In practical scenarios, it is often necessary to segment individual plumes from an image and estimate the emission rates separately.
@@
Since the introduction of AlexNet in 2012, deep learning has witnessed rapid development and yielded remarkable results in various fields of remote sensing application[JIA202314][JIANG202330][HE202476]. The U-net deep learning architecture was initially designed for semantic segmentation tasks[ronneberger2015u][POLEWSKI2021297], but also exhibits enormous potential for regression[10.1145/3429309.3429328]. We regard methane concentration inversion as pixel-by-pixel regression, so U-net can be used to complete the task. Additionally, Mask R-CNN has proven to be highly effective in a wide range of instance segmentation tasks [SANIMOHAMMED2022100024] and should also be applicable to Methane plume segmentation. Moreover, ResNet addresses the issue of gradient vanishing, enabling deeper networks to improve feature extraction capabilities[he2016deep][shafiq2022deep]. Several earlier studies on emission rate estimation using shallow convolutional neural networks (CNN) have demonstrated feasibility, albeit with limited accuracy[9027926][guanter2021mapping]. It is reasonable to apply ResNet to improve the emission rate estimation. And our goal is to leverage physically simulated plumes as training data for multi-task deep learning techniques to simultaneously perform methane concentration inversion, methane plume segmentation, and flux rate estimation in a unified approach.
@@
In this article, we start by describing the foundational principles of the algorithms utilized in our study, including large-eddy simulation, radiation simulation, the methane plume concentration inversion algorithm, the methane plume segmentation algorithm (instance segmentation), and the plume emission rate estimation algorithm. Following this, we provide a detailed account of our experimental setup, which encompasses the research area, the collected data, and the procedures used to build our methane monitoring dataset. Lastly, we evaluate the performance of deep learning algorithms against traditional ones for the tasks of methane plume concentration inversion, methane plume segmentation (instance segmentation) and plume emission rate estimation. Furthermore, we present the two multi-task learning networks, MTL-01 and MTL-02, that we developed for these tasks. The major contributions of this paper are as follows:
@@
1. A physics-informed deep learning scheme is proposed for spaceborne retrieval of methane emissions based on data simulation using Large Eddy Simulation (LES), Radiative Transfer Equation (RTE), and data augmentation techniques.
@@
2. An instance segmentation algorithm is applied to solve the problem of isolating and identifying methane emission sources by plume segmentation.
@@
3. We show that the serialization of multiple sub-tasks for methane emission monitoring leads to additional error accumulation, and designed a multi-task learning model to suppress the errors.
@@ II-A Simulation
The simulation operation consists of two parts: simulation of methane plume emission and simulation of EnMAP-like satellite images with methane plume signal. For methane plume simulation, we utilized Large Eddy Simulation (LES), which is a computational fluid dynamics (CFD) method used to simulate turbulent phenomena in fluid flow[sagaut2002large]. It combines the advantages of Direct Numerical Simulation (DNS) and Reynolds-averaged Navier-Stokes equations (RANS) methods, aiming to accurately predict the statistical characteristics of turbulent flow[pope2001turbulent][domaradzki1997subgrid]. The core of LES lies in filtering the flow field into large- and small-scale vortices through a filtering operation. This can be achieved by applying a low-pass filter to remove high-frequency small-scale vortices while retaining the low-frequency large-scale vortices[domaradzki1997subgrid]. The filtered flow field equations form the basis of LES. The governing equations for LES are based on Reynolds-averaged Navier-Stokes equations (RANS)[germano1991dynamic], which are solved to simulate motion and turbulent effects in the fluid [10.1115/1.3124648]:
@@
By combining these two steps, we can simulate methane plumes of various shapes, sizes, and concentrations and inject them into EnMAP data to generate simulated hyperspectral satellite images, which contain methane plume signals overlaid with a realistic scene background. This provides a solid foundation for our research and allows us to validate the performance and reliability of our methodology in real-world scenarios.
@@ II-B Methane Concentration Inversion
Methane concentration inversion typically refers to a technique or process utilized to retrodict (or estimate) the concentration of methane on the earth’s surface or in the atmosphere from satellite images or other remote sensing data. This process involves the application of mathematical models and algorithms to transform satellite imagery into quantitative assessments of methane concentrations.
@@
Methane concentration inversion in the West New Territories (WENT) region was performed using PRISMA data and the mag1c method. The results of this inversion were evaluated against a deep learning model to validate its generalization capabilities. Considering the intricate distribution of land and sea areas in Hong Kong, as well as the challenging rugged terrain and dense urban construction, it can be expected that the surface conditions in the region significantly influence the results of methane concentration inversion. To address this, we use a masking method based on K-means to run Mag1c[funk2001clustering].
In the realm of deep learning technologies, U-net is a prevalent and easily implementable architecture, frequently used for image semantic segmentation tasks[ronneberger2015u][POLEWSKI2021297][zang2021land]. Our rationale for selecting U-net as the model for methane concentration inversion comes from the following points: 1. We aim to achieve inversion of the methane concentration using a network that is as simplistic as possible. 2. Within the U-net framework, semantic segmentation is realized in a regression form, making it conveniently adaptable to the task of methane concentration inversion. 3. U-net can easily produce output images with the same length and width as the input image. Typically, the input and output channels of the U-net are of equal size. However, to accommodate the requirements for the inversion of methane concentration, an additional layer of $1*1$ convolution is added after the U-net encoding-decoding structure[lin2013network], and the loss function is replaced for the regression task. The loss functions used include the mean squared error (MSE) and $SmoothL_{1}$:
@@
Before starting the training phase, various preprocessing techniques were applied to both labels and images. In particular, we computed the mean and standard deviation of all non-zero pixel values in the labels to achieve normalization. Referring to the pre-processing methods of the DOAS algorithm [wagner2004max][platt2008differential], the subsequent formula was utilized to preprocess the images:
@@
where $I_{before}$ denotes the image prior to preprocessing, and $I_{after}$ denotes the same image post preprocessing. Both EnMAP and PRISMA have pixel values within the range of 0-10 after taking the logarithm, and dividing these values by 10 normalizes all preprocessed pixel values to fall within the range of 0-1. This technique helps to generalize the model across various sensors. The network is optimized using two loss functions, MSE and SmoothL1, starting with an initial learning rate of 0.001. The learning rate is decreased by 75% every 10 epochs, over a total of 200 epochs.
@@ II-C Methane Plume Segmentation
Instance segmentation is a classical task in computer vision, and methane plume segmentation is essentially an instance segmentation task. Previous research work on methane plume segmentation is limited, and many studies are based on subjective segmentation criteria[gorrono2023understanding]. Some studies suggested using traditional algorithms for the segmentation of methane plumes, with contour tracking algorithms considered a good choice due to their efficiency and stability. On the other hand, there are many deep learning-based instance segmentation methods, among which Mask R-CNN is the most classical and widespread one[he2017mask]. It is a two-step instance segmentation algorithm that has proven to achieve high performance in a large number of datasets for a wide range of applications. Although new instance segmentation algorithms are continuing to emerge, some of the algorithms sacrifice segmentation accuracy for real-time operation[terven2023comprehensive], which is not necessarily required for our case. Other algorithms may achieve higher segmentation accuracy[wang2020solov2], but require an increasingly complex optimization process during training, which is not cost-effective. Therefore, Mask R-CNN is chosen as the deep learning algorithm for methane plume segmentation and active contour algorithms as the traditional counterpart. Typically, Mask R-CNN takes three image channels as inputs, but we modified the input layer to allow flexible alteration between 1 or 41 input channels to accommodate different segmentation requirements[foote2020fast], since the methane distribution map has only one band and EnMAP contains 41 bands after excluding the water vapor and carbon dioxide sensitive bands. The original loss function of Mask R-CNN is used:
@@
where $L$ is the total loss function, $L_{cls}$ is the category loss function, which is based on the cross-entropy loss, $L_{box}$ is the bounding box loss function(As Mask R-CNN is a two-step instance segmentation model,its loss function inevitably includes the loss of object detection task), which is based on the $SmoothL_{1}$ loss, and $L_{mask}$ is the mask loss function, which typically uses the binary cross entropy:
@@
where $n$ is the number of bounding boxes considered in the current batch, $x_{i}$ represents a parameter set (including the center coordinates, width and height of the box) of the $i$-th predicted bounding box, $\hat{x}_{i}$ is the corresponding ground truth parameter for the $i$-th bounding box. During the training process, we use the minimum enclosing rectangle of the ground truth masks as the bounding box (for the predicted and ground truth bounding box).
$\text{Smooth}_{L1}(x)=\begin{cases}0.5x^{2}&\text{if }|x|<1\\ |x|-0.5&\text{if }|x|\geq 1\end{cases}$ (10)
where $x$ is the difference between the predicted value and the true value.
$L_{mask}=-\text{IoU}(b,\hat{b})+\frac{\rho^{2}(b,\hat{b})}{c^{2}}$ (11)
where $b$ and $\hat{b}$ respectively denote the ground truth binary mask and the predicted binary mask, $\text{IoU}(b,\hat{b})$ indicates the intersection over the union between these two binary masks, $\rho^{2}(b,\hat{b})$ is the Euclidean distance between the center points of two binary masks, and $c$ is a hyperparameter typically set to 10.
II-D Emission Rate Estimation
There are two scenarios for estimating emission rates. The first scenario involves estimating the emission rates using a supervised deep model with training samples collected specifically for this purpose. The goal here is to evaluate the performance of the estimation algorithm under ideal conditions. The second scenario refers to the estimation of emission rates based on plume segmentation results (MTL-01), where the objective is to evaluate and improve the accuracy of plume instance segmentation.
@@ II-D Emission Rate Estimation
ResNet is a deep learning method that uses residual blocks, which enables the neural network to learn the residual map by effectively avoiding issues such as gradient vanishing and explosion[he2016deep]. ResNet is widely applied in various application fields, including image recognition, object detection, and speech processing, yielding impressive results[veit2016residual][shafiq2022deep]. However, in this study, we pay particular attention to the application of ResNet in regression tasks, specifically for the complex and important task of emission rate estimation. Since the emission process is often affected by a variety of environmental factors and source dynamics, this makes emission rate estimation a highly nonlinear and complex problem. To meet this challenge, we chose ResNet-50 as our base network architecture for the following reasons: 1. Powerful feature extraction capabilities: ResNet-50 is able to learn and extract rich hierarchical features in an image through deep structure and residual connection, which is essential to capture subtle changes in methane emission from images. 2. Suitable for regression tasks: although ResNet-50 was originally designed for classification tasks, it can be easily adapted to regression tasks, such as emission rate estimation, by adjusting the output layer of the network (replacing the softmax layer with a linear layer).
We trained the ResNet-50 network with samples to perform the flux-rate estimation task. Subsequently, ResNet-50 was used as a pretrained network to validate the rationale of the plume segmentation results and assist in optimizing the instance segmentation network.